Design a small gas turbine engine to produce 130kW of net power. Use an air-standard analysis (const. gas prop.) and assume air enters the compressor at 100kPa and 20^\circ C . the compressor pressure ratio is 10, the maximum cycle temperature is 750^\circ C , and the cold air stream leaves the regenerator 15^\circ C cooler than the hot air stream at the inlet of the regenerator. Assume a compressor isentropic efficiency of 87% and a turbine isentropic efficiency of 95%. Determine the rates of heat addition and rejection required to produce 130kW. cp=1.005 kJ/kg
Answer #1

TA 7500 26_3 2 1-2 - Compression 2-3 → Gnst.prepuye heat add 3-4 = Sentropic Capanion 4-1 = const Pressure A 200 PEC 7QR heat1:-) та е (10) 1, 293 Т. 565-696 (+ в ка» 4] . 14-) (lb) гч o 23. 4 Т - . 529. 86 ), Gon: Lлитерс ара сте се мрти - 8*lb - еGiven What Wp - c = 130kW Where Wp = mcp [T3 - Ty Wes Mice [T- Whene mos mais facio rate 130 = rcp ( 13 - Tu - 72 +T BD. = X